OpenStudy (anonymous):

MEDAL Simplify. (3a – 4)2 A. 9a2 – 16 B. 9a2 – 24a + 16 C. 6a2– 24a + 8 D. 92a – 7 + 16

OpenStudy (anonymous):

@Nnesha @dtan5457

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use the identity \[(x-y)^2 = x^2-2xy+y^2\]

OpenStudy (anonymous):

ok i really suck at this, and i dont really understand what we have to do

ganeshie8 (ganeshie8):

easy, \[(3a – 4)^2=(3a)^2-2(3a)(4)+4^2\] simplify

OpenStudy (anonymous):

alright so, 9a^2- 24a+16

ganeshie8 (ganeshie8):

Excellent!